Journalist Report – November 10th

SOL: 12
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: No
Currently charging: Yes
Opportunity rover used: No
Currently charging: Yes
Curiosity rover used: No
Currently charging: Yes
Perseverance rover used: No
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: unknown, but the difference between the used water and the added water left us with 52 gallons more than yesterday
Water (static tank): 265 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: no
Summary of internet: No comments
Summary of suits and radios: Suit number 9 was not blowing as much air as usual when tested before the EVA. That suit was not used in the EVA but rather left in the base to test it further. After 15 min with the fans turned on, the voltage measured 2.3 V, so the battery was replaced and the old one marked as faulty. The new battery was measured before and after charge with nominal results.
Summary of GreenHab operations: watered, harvested
WATER USE: 2.5 gallons
Heater: On
Supplemental light: Off
Harvest: 20 g of radish
Summary of ScienceDome operations: Not used today
Dual split: Off
Summary of RAM operations: Not used today
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site